# MUG-V 10B: High-efficiency Training Pipeline for Large Video Generation Models

### v1

@@ 1 Introduction
In this work, we focus on developing a high-efficiency training framework for diffusion transformers (DiT) and training a large-scale video generation model. Video generation is among the most challenging forms of visual content synthesis: relative to image generation, it must preserve static content fidelity while learning diverse motion dynamics; relative to 3D generation, it should not only implicitly capture object-level 3D structure but also model inter-object interactions and physical regularities wang2025survey ; lu2024handrefiner . Meanwhile, training large-scale video generation models also demands addressing three core challenges: cross-modal text-video alignment, extra long visual token sequences and complex spatiotemporal patterns wen2025analysis ; zeng2024dawn ; kaddour2023challenges ; zhang2024vast ; chen2025temporal .
@@ 2.1 Scalable Video Processing Pipeline
Caption generation. High-fidelity captions drive both prompt-following and training convergence. We first finetune a Qwen2-VL-72B Qwen2VL ; Qwen-VL captioner on public datasets plus internally labelled clips, optimising for descriptions that cover objects, appearance, motion, and background context. The capability is then distilled into a Qwen2-VL-7B model Qwen2VL ; Qwen-VL , striking a balance between accuracy and inference throughput for large-scale captioning.
@@ 4.2.3 Post-training and Alignment
After the multi-stage pre-training described above, the validation loss plateaued and began to oscillate, the model’s outputs exhibited two persistent failure modes: (i) fine-grained artefacts, especially in articulated regions such as human hands, (ii) violations of basic physical plausibility (e.g., interpenetration and distortions). To further improve generative quality we adopted two post-training approaches: annealed supervised fine-tuning (SFT) with post-EMA, and preference-based optimization.
@@ 7.2 Video Generation Models
A dominant family today uses DiT-based generators operating in a latent video space: a video VAE provides compact spatiotemporal latents, while a DiT (with factorized or windowed spatiotemporal attention) performs conditional generation under text, image, or control signals zhou2022magicvideo ; chen2023videocrafter1 . Advances include stronger conditioning (pose, depth, camera paths, audio), longer context handling (memory-efficient attention, sliding windows), and faster sampling via consistency or flow matching objectives. Large proprietary systems, e.g., Sora demonstrate long-duration, high-resolution generation with improved physical plausibility through large-scale training, aggressive latent compression, and optimized inference brooks2024video ; deepmind_veo3_models_2025 ; wan2025wan ; kong2024hunyuanvideo ; gao2025seedance ; ma2025step ; chen2025goku ; team2025yan .
In parallel, autoregressive (AR) based approaches tokenize videos with vector-quantized encoders and model generation as next-token prediction across spatiotemporal tokens yan2021videogpt ; yu2023magvit . These models integrate naturally with multimodal LLMs and show strengths in long-horizon structure and discrete controllability, but often trade off visual fidelity and suffer from compression artifacts. Hybrid systems combine AR planning (for structure and semantics) with diffusion/flow decoders (for photorealism), narrowing this gap kondratyuk2023videopoet ; chen2024videocrafter2 .
@@ A.2 More Details of Preference Optimisation
In addition to direct preference optimization (DPO) and KTO with human-labeled data, we introduce Real Data Preference Optimization (RDPO) qian2025rdpo . By applying reverse sampling on real video data, we observe that the flow sampling paths derived from these reverse samples are statistically superior to those obtained from randomly initialized noise and its associated flow trajectories. This property allows RDPO to automatically construct preference pairs without the need for manual annotation. Furthermore, a multi-stage iterative training schedule is employed to progressively improve the generator’s performance.

### v2

@@ 1 Introduction
In this work, we focus on developing a high-efficiency training framework for diffusion transformers (DiT) and training a large-scale video generation model. Video generation is among the most challenging forms of visual content synthesis: relative to image generation, it must preserve static content fidelity while learning diverse motion dynamics; relative to 3D generation, it should not only implicitly capture object-level 3D structure but also model inter-object interactions and physical regularities wang2025survey ; lu2024handrefiner ; wiedemer2025video . Meanwhile, training large-scale video generation models also demands addressing three core challenges: cross-modal text-video alignment, extra long visual token sequences and complex spatiotemporal patterns wen2025analysis ; zeng2024dawn ; kaddour2023challenges ; zhang2024vast ; chen2025temporal .
@@ 2.1 Scalable Video Processing Pipeline
Caption generation. High-fidelity captions are crucial for both prompt adherence and stable training convergence, which currently rely heavily on advanced video understanding models Qwen-VL ; wu2025number . We first finetune a Qwen2-VL-72B Qwen2VL captioner on public datasets plus internally labelled clips, optimising for descriptions that cover objects, appearance, motion, and background context. The capability is then distilled into a Qwen2-VL-7B model Qwen2VL , striking a balance between accuracy and inference throughput for large-scale captioning.
@@ 4.2.3 Post-training and Alignment
After the multi-stage pre-training, the validation loss plateaued and began to oscillate, the model’s outputs exhibited two persistent failure modes: (i) fine-grained artefacts, especially in articulated regions such as human hands, (ii) violations of basic physical plausibility (e.g., interpenetration and distortions). To further improve generative quality we adopted two post-training approaches: annealed supervised fine-tuning (SFT) with post-EMA, and preference-based optimization karras2024analyzing ; rafailov2023direct ; peng2025lmm .
@@ 7.2 Video Generation Models
A dominant family today uses DiT-based generators operating in a latent video space: a video VAE provides compact spatiotemporal latents, while a DiT (with factorized or windowed spatiotemporal attention) performs conditional generation under text, image, or control signals zhou2022magicvideo ; chen2023videocrafter1 . Advances include stronger conditioning (pose, depth, camera paths, audio), longer context handling (memory-efficient attention, sliding windows), and faster sampling via consistency or flow matching objectives. Large proprietary systems, e.g., Sora sora_openai_2024 , Sora2 sora2_openai_2025 , Veo3 deepmind_veo3_models_2025 demonstrate long-duration, high-resolution generation with improved physical plausibility through large-scale training, aggressive latent compression, and optimized inference brooks2024video ; wan2025wan ; kong2024hunyuanvideo ; gao2025seedance ; ma2025step ; chen2025goku ; team2025yan .
In parallel, autoregressive (AR) based approaches tokenize videos with vector-quantized encoders and model generation as next-token prediction across spatiotemporal tokens yan2021videogpt ; yu2023magvit . These models integrate naturally with multimodal LLMs and show strengths in long-horizon structure and discrete controllability, but often trade off visual fidelity and suffer from compression artifacts. Hybrid systems combine AR planning (for structure and semantics) with diffusion/flow decoders (for photorealism), narrowing this gap kondratyuk2023videopoet ; chen2024videocrafter2 ; wang2025ovis ; xie2025show ; agarwal2025cosmos .
@@ A.2 More Details of Preference Optimisation
In addition to direct preference optimization (DPO) and KTO with human-labeled data, we introduce Real Data Preference Optimization (RDPO) qian2025rdpo . By applying reverse sampling on real video data, we observe that the flow sampling paths derived from these reverse samples are statistically superior to those obtained from randomly initialized noise and its associated flow trajectories. This property allows RDPO to automatically construct preference pairs without the need for manual annotation. Furthermore, a multi-stage iterative training schedule is employed to progressively improve the generator’s performance. During post-training, we observe that well-tuned reinforcement learning can rapidly strengthen the model’s generative capabilities along specific dimensions, however, striking an effective trade-off between supervised fine-tuning (SFT) and RL remains a challenging open problem wu2025generalization .